# Red Teaming AI Red Teaming

### v1

@@ Abstract
Red teaming has evolved from its origins in military applications to become a widely adopted methodology in cybersecurity and AI. In this paper, we take a critical look at the practice of AI red teaming. We argue that despite its current popularity in AI governance, there exists a significant gap between red teaming’s original intent as a critical thinking exercise and its narrow focus on discovering model-level flaws in the context of generative AI. Current AI red teaming efforts focus predominantly on individual model vulnerabilities while overlooking the broader sociotechnical systems and emergent behaviors that arise from complex interactions between models, users, and environments. To address this deficiency, we propose a comprehensive framework operationalizing red teaming in AI systems at two levels: macro-level system red teaming spanning the entire AI development lifecycle, and micro-level model red teaming. Drawing on cybersecurity experience and systems theory, we further propose a set of recommendations. In these, we emphasize that effective AI red teaming requires multifunctional teams that examine emergent risks, systemic vulnerabilities, and the interplay between technical and social factors.
@@ 1 Introduction
In recent past, the term red teaming has gained significant attention in diverse conversations around AI as a potential solution to find and address security, safety, and reliability concerns in generative AI (genAI) systems. When ChatGPT came out in November 2022, not many people in the field of AI except those who had previously been in the military or cybersecurity knew about red teaming. Less than a year later, the largest AI red teaming exercise ever was organized at DEF CON, a leading hacker conference [7, 3]. Fast forward to today, where we are in the middle of a (supposed) genAI revolution, AI red teaming dominates conversations in academia [20, 21], think tank advice [27, 45], and popular media articles [6, 4, 35]. The notion of red teaming is being touted as a panacea to many of the problems AI might have, and its application to genAI systems has been and continues to be encouraged as a matter of practice. While that is not necessarily a bad thing—because red teaming is definitely a part of solving the AI security and safety problem—it is only a small part of a more complete set of tools that should be used.
@@
There is a critical disconnect between the rich technical foundation and industry best practices in cybersecurity and the current approach to AI red teaming. Despite decades of research and education in security, today’s AI red teaming efforts remain narrowly focused on technical vulnerabilities in genAI models (or model-driven general purpose systems like ChatGPT and Claude). This missing link negates the tremendous benefits that are unrealized in bridging communities of practice in AI and cybersecurity, which could ultimately lead to us finding ways to achieve robust and effective AI governance and AI solutions that adhere more closely to the espoused principles of responsible AI.
To effectively implement red teaming in a sociotechnical system like the AI development lifecycle, we need to expand our perspective beyond the narrow technical focus that currently dominates the discourse. Red teaming originated as a strategic thinking exercise in which a designated team not only simulates adversarial actions, but challenges assumptions and identifies blind spots as part of a well-defined project greenlighting process. The current approach of testing something that has already been built, as many AI companies are doing [2, 12], misses the point of pre-emptive and proactive critical analysis that red teaming was originally designed to provide.
@@ 1.1 What Red Teaming is and is not
Red teaming is fundamentally a team exercise [53]. It should bring together multi-functional and cross-functional teams around the common goal of ultimately improving a product or user experience. The means of this exercise are to explore hypotheticals centered around two primary questions: (1) Does the system work for their interests? (2) Could it work against their interests? The multidisciplinary approach acknowledges that systems built for a specific purpose exist within complex sociotechnical environments where technical, social, and organizational factors interact. Putting things in context, in AI red teaming technical teams are best suited to investigate model vulnerabilities. On the other hand, policy experts help identify regulatory conflicts, ethicists can surface value alignment issues, and domain specialists can evaluate real-world impact scenarios.
@@ 2.1 Military Origins
Red teaming emerged from the tactical war games of the Prussian military in the early 19th century, evolving through Cold War simulations into today’s formalized methodology for challenging conventional wisdom and identifying strategic and exploitable hazards. The Prussian army adopted Kriegsspiel (literally “wargame” in German) in 1812, a tabletop simulation developed by Lieutenant Georg Leopold von Reisswitz and his son, where blue pieces represented the Prussian forces and red represented the enemy [28, 22]. This color-coding established the concept of “red team” that persists to this day.
In the United States (US), the practice of red teaming took shape during the Cold War, when the RAND Corporation conducted military simulations for the US government. In these exercises, “red team” and the color red were used to represent the Soviet Union, while “blue team” and blue represented the United States [51]. Following the intelligence failures that led to the September 11 attacks in 2001, the US Department of Defense established formal red team units to prevent similar catastrophic oversights in future. The 9/11 Commission identified a “failure to connect the dots” as a primary cause of the intelligence breakdown. This led to systematic changes to prevent groupthink and foster alternative analysis [38]. As a result, the Pentagon created specialized training centers such as the University of Foreign Military and Cultural Studies at Fort Leavenworth to institutionalize red teaming methodologies. The US Army’s University of Foreign Military and Cultural Studies at Fort Leavenworth, created after intelligence failures in Iraq, developed the modern framework for red teaming that transformed an ad hoc practice into a systematic methodology for critical analysis [14]. This program teaches military officers and government officials techniques to challenge assumptions, consider alternative perspectives, and introduce contrarian thinking into planning processes.
The concept of the “10th man” from Israeli military doctrine, popularized in the film “World War Z”, illustrates another approach to institutionalized contrarian thinking. When everyone agrees on a particular outcome, it is the designated 10th man’s responsibility to disagree and explore alternative scenarios. This concept reportedly developed after intelligence failures during the 1973 Yom Kippur War, when analysts unanimously agreed that Arab troop movements weren’t a threat [28]. In reality, while Israeli intelligence did establish a unit called Ipcha Mistabra (“the opposite side”) to challenge prevailing assumptions after the Yom Kippur War, the specific “10th man” concept as portrayed in the film is somewhat fictionalized but based on real adversarial thinking practices in military intelligence [51].
@@ 2.2 Adoption in Cybersecurity
The National Security Agency (NSA) first recognized the need for proactive cybersecurity measures in the 1980s, and pioneered the concept of “red teams” tasked with assessing the security of classified systems [46]. These early efforts involved independent evaluators simulating potential attackers and identifying weaknesses that required remediation.
@@
As digital threats evolved in the 1990s, so did cybersecurity red teaming. The term “tiger team” was initially used to describe specialized groups that performed many of the same functions as modern red teams [16]. These groups of expert professionals were hired to take on particular challenges against the security posture of organizations.
@@
Following the 9/11 attacks, cybersecurity red teaming gained significant momentum as organizations recognized the need for more comprehensive security testing. The Central Intelligence Agency created a new unit called “Red Cell”, and red teaming became increasingly common in various government agencies to model responses to asymmetric threats, including cyber attacks [51]. This period marked the transition from isolated pentesting to more holistic security assessments that incorporated physical security, social engineering, and other non-technical aspects.
@@
Modern cybersecurity red teaming encompasses several key methodologies working in concert: technical assessments that test digital defenses through vulnerability scanning, exploitation, and lateral movement [19]; physical security testing that evaluates access controls for facilities [41]; social engineering that targets the human element through phishing and impersonation [48]; and extended red team operations designed to achieve specific objectives while testing detection and response capabilities [39]. These approaches are codified in frameworks such as NIST Special Publication 800-53, which includes specific controls for red team exercises designed to “simulate attempts by adversaries to compromise organizational information systems” and “provide comprehensive assessments that reflect real-world conditions” [11].
@@
The field continues to evolve with several advanced approaches. Continuous Automated Red Teaming (CART) uses automation to assess security posture in real-time rather than through periodic manual assessments [19]. Adversary Emulation models tactics after specific threat actors that might target the organization, guided by frameworks like MITRE ATT&CK [8]. Purple Teaming fosters collaboration between red and blue teams to identify vulnerabilities and improve response strategies [13]. Integrated IT-OT Assessments expand scope to include industrial control systems and critical infrastructure [17]. AI-Enhanced Red Teaming that incorporates AI to improve the effectiveness of assessments [23]. And finally, specialized services from organizations like CISA provide comprehensive evaluations for critical infrastructure sectors and government agencies [8].
@@ 2.3 Red Teaming AI
A working definition of the very new concept of AI red teaming is that it is structured testing to identify flaws and vulnerabilities in AI systems, typically conducted in controlled environments with developer collaboration. For Large Language Models (LLMs) specifically, red teaming is defined as “a process where participants interact with the LLM under test to help uncover incorrect or harmful behaviors” [40]. LLM development companies have implemented various approaches to red teaming, ranging from comprehensive security assessments to narrower evaluations focused on specific genAI features [5]. In the last two years, such approaches have become recognized as essential for assessing the risks of AI models and systems [1].
Despite its growing popularity, researchers have identified significant challenges with current AI red teaming practices. The field lacks consensus on the scope, structure, and assessment criteria for AI red teaming [15], raising concerns that red teaming may sometimes function more as performative activity than substantive risk mitigation. Many current approaches focus too narrowly on the models themselves, neglecting how vulnerabilities might manifest in production systems where AI models are a part of broader systems [5]. Additionally, most AI experts fail to consider insider risks [26], and most testing processes remain limited to English-language evaluations [40].
The process of red teaming itself presents challenges for participants. They can experience negative psychological impacts from interaction with harmful content and adversarial thinking, which can lead to decreased productivity or psychological harm [1]. Many red teamers also lack training in crucial disciplines outside of their technical expertise, with most employee red teamers having limited proficiency in linguistic, sociocultural, historical, legal, or ethical domains [18].
As the field matures, researchers increasingly recognize that red teaming alone cannot solve all the challenges in AI risk assessment [1]. More effective approaches such as violet teaming are emerging that integrate complementary methods, combining red teaming’s risk identification capabilities with the solution-focused responses of blue teaming [47]. Future success will likely depend on embracing greater diversity and focusing more on practical red teaming efforts that address real-world attacks, which are often less sophisticated than those presented in academic papers [37].
@@ 2.4 What’s Missing?
As such, the state-of-the-art of AI red teaming overindexes on model-specific behaviors rather than how these models work within AI applications, interact with broader social contexts, and affect the external world through their outputs. Current AI red teaming efforts tend to focus narrowly on technical vulnerabilities while overlooking broader sociotechnical considerations. The public interest dimension of red teaming remains underdeveloped as well [45].
To mitigate these flaws, AI red teams need to embrace cybersecurity’s decades of experience with security testing and reporting. A large proportion of the testing done in the name of red teaming AI systems can be structured and automated through frameworks like MITRE ATT&CK for understanding adversarial behaviors, prioritizing vulnerabilities and coordinating defensive responses [31], and established practices for continuous monitoring, automated testing, and incident response [32]. In comparison, true red teaming requires “an alchemist mindset” that extends beyond purely technical approaches [20]. Successful cyber red team engagements typically involve creating diverse and highly realistic scenarios that produce actionable insights to aid proactive mitigation [52]—principles that are often missing from AI red teaming.
@@ 3.1 Macro-level (System) Red Teaming
Just like technical debt in ML systems can arise from system components other than code [43], ML (and AI) failures can stem from decisions made long before the first line of code is written (Figure 1).
Figure 1: Technical and nontechnical components of a ML system, with those with potential trust considerations marked by blue circles. Reproduced with permission from Pruksachatkun et al. [36].
@@ 3.1.1 Inception
The inception stage is critical, given this is where stakeholders first envision an AI solution to address a particular challenge. During this stage, the job of macro-level red teaming is to challenge fundamental assumptions about problem framing and solution appropriateness. Red teams at this stage should ask probing questions: Is AI actually necessary to solve this problem, or are we succumbing to “AI solutionism”? What are the stakeholder motivations driving this project, and how might different constituencies be affected? What adjacent problems might we inadvertently create? How might adversaries exploit or manipulate this system once deployed? [49]
@@
Critically, they should help everyone be on the same page to answer the question “What does good look like?” To this end, stakeholders must articulate not just what they want the system to do, but what they absolutely cannot allow it to do. This negative space definition often reveals assumptions and constraints that were not initially apparent [45].

### v2

@@ Abstract
Red teaming has evolved from its origins in military applications to become a widely adopted methodology in cybersecurity and AI. In this paper, we take a critical look at the practice of AI red teaming. We argue that despite its current popularity in AI governance, there exists a significant gap between red teaming’s original intent as a critical thinking exercise and its narrow focus on discovering model-level flaws in the context of generative AI. Current AI red teaming efforts focus predominantly on individual model vulnerabilities while overlooking the broader sociotechnical systems and emergent behaviors that arise from complex interactions between models, users, and environments. To address this deficiency, we propose a comprehensive framework operationalizing red teaming in AI systems at two levels: macro-level system red teaming spanning the entire AI development lifecycle, and micro-level model red teaming. Drawing on cybersecurity experience and systems theory, we further propose a set of six recommendations. In these, we emphasize that effective AI red teaming requires multifunctional teams that examine emergent risks, systemic vulnerabilities, and the interplay between technical and social factors.
keywords:
Red teaming, AI red teaming, AI security, AI vulnerability
@@ 1 Introduction
In recent past, the term red teaming has gained significant attention in diverse conversations around AI as a potential solution to find and address security, safety, and reliability concerns in generative AI (genAI) systems. When ChatGPT came out in November 2022, red teaming was a relatively unknown term for AI researchers and practitioners without a background in military or cybersecurity. Less than a year later, the largest AI red teaming exercise ever was organized at DEF CON, a leading hacker conference (Cattell et al., 2023; Bajak, 2023). Fast forward to today, where we are in the middle of a (supposed) genAI revolution, AI red teaming dominates conversations in academia (Inie et al., 2025; Ji and Shea-Blymyer, 2024), think tank advice (Metcalf and Singh, 2024; Singh et al., 2025), and popular media articles (Burt, 2024; Barker, 2025; Patwa, 2025). The notion of red teaming is being touted as a panacea to many of the problems AI might have, and its application to genAI systems has been and continues to be encouraged as a matter of practice. While that is not necessarily a bad thing—because red teaming is definitely a part of solving the AI security and safety problem—it is only a small part of a more complete set of tools that should be used.
@@
There is a critical disconnect between the rich technical foundation and industry best practices in cybersecurity and the current approach to AI red teaming. Despite decades of research and education in security, today’s AI red teaming efforts remain narrowly focused on technical vulnerabilities in genAI models (or model-driven general-purpose systems like ChatGPT and Claude). This missing link negates the tremendous benefits that are not realized in bridging communities of practice in AI and cybersecurity, which could ultimately lead us to finding ways to achieve robust and effective AI governance and AI solutions that adhere more closely to the espoused principles of responsible AI.
To effectively implement red teaming in a sociotechnical system like the AI development lifecycle, we need to expand our perspective beyond the narrow technical focus that currently dominates the discourse. Red teaming originated as a strategic thinking exercise in which a designated team not only simulates adversarial actions, but challenges assumptions and identifies blind spots as part of a well-defined project greenlighting process. The current approach of testing something that has already been built, as many AI companies are doing (Anthropic, 2024; Daws, 2024), misses the point of pre-emptive and proactive critical analysis that red teaming was originally designed to provide.
@@ 1.1 What Red Teaming is and is not
Red teaming is fundamentally a team exercise Zenko (2015). It should bring together multi-functional and cross-functional teams around the common goal of ultimately improving a product or user experience. The means of this exercise are to explore hypotheticals centered around two primary questions: (1) Does the system work for their interests? (2) Could it work against their interests? The multidisciplinary approach acknowledges that systems built for a specific purpose exist within complex sociotechnical environments where technical, social, and organizational factors interact. Putting things in context, in AI red teaming technical teams are best suited to investigate model vulnerabilities. On the other hand, policy experts help identify regulatory conflicts, ethicists can surface value alignment issues, and domain specialists can evaluate real-world impact scenarios.
@@ 1.2 The Way Forward
The remainder of this paper proceeds as follows. Section 2 traces the evolution of red teaming from military war games through cybersecurity practices to current AI applications, highlighting the progressive narrowing of scope. Section 3 presents our two-level framework, distinguishing macro-level system red teaming from micro-level model testing, and provides concrete examples of each. Section 4 presents a few recommendations to consider AI red teaming from a systems perspective, and Section 5 concludes the paper.
@@ 2.1 Military Origins
Red teaming emerged from the tactical war games of the Prussian military in the early 19th century, evolving through Cold War simulations into today’s formalized methodology for challenging conventional wisdom and identifying strategic and exploitable hazards. The Prussian army adopted Kriegsspiel (literally “wargame” in German) in 1812, a tabletop simulation developed by Lieutenant Georg Leopold von Reisswitz and his son, where blue pieces represented the Prussian forces and red represented the enemy (Meyer, 2025; Kay, 2020). This color-coding established the concept of “red team” that persists to this day.
In the United States (US), the practice of red teaming took shape during the Cold War, when the RAND Corporation conducted military simulations for the US government. In these exercises, “red team” and the color red were used to represent the Soviet Union, while “blue team” and blue represented the United States (Wikipedia, 2025). Following the intelligence failures that led to the September 11 attacks in 2001, the US Department of Defense established formal red team units to prevent similar catastrophic oversights in future. The 9/11 Commission identified a “failure to connect the dots” as a primary cause of the intelligence breakdown. This led to systematic changes to prevent groupthink and foster alternative analysis (Red Team Thinking, 2021). As a result, the Pentagon created specialized training centers such as the University of Foreign Military and Cultural Studies at Fort Leavenworth to institutionalize red teaming methodologies. The US Army’s University of Foreign Military and Cultural Studies at Fort Leavenworth, created after intelligence failures in Iraq, developed the modern framework for red teaming that transformed an ad hoc practice into a systematic methodology for critical analysis (Dooley, 2017). This program teaches military officers and government officials techniques to challenge assumptions, consider alternative perspectives, and introduce contrarian thinking into planning processes.
The concept of the “10th man” from Israeli military doctrine, popularized in the film “World War Z”, illustrates another approach to institutionalized contrarian thinking. When everyone agrees on a particular outcome, it is the designated 10th man’s responsibility to disagree and explore alternative scenarios. This concept reportedly developed after intelligence failures during the 1973 Yom Kippur War, when analysts unanimously agreed that Arab troop movements weren’t a threat (Meyer, 2025). In reality, while Israeli intelligence did establish a unit called Ipcha Mistabra (“the opposite side”) to challenge prevailing assumptions after the Yom Kippur War, the specific “10th man” concept as portrayed in the film is somewhat fictionalized but based on real adversarial thinking practices in military intelligence (Wikipedia, 2025).
@@ 2.2 Adoption in Cybersecurity
The National Security Agency (NSA) first recognized the need for proactive cybersecurity measures in the 1980s, and pioneered the concept of “red teams” tasked with assessing the security of classified systems (TechRound, 2023). These early efforts involved independent evaluators simulating potential attackers and identifying weaknesses that required remediation.
@@
As digital threats evolved in the 1990s, so did cybersecurity red teaming. The term “tiger team” was initially used to describe specialized groups that performed many of the same functions as modern red teams (Firch, 2024). These groups of expert professionals were hired to take on particular challenges against the security posture of organizations.
@@
Following the 9/11 attacks, cybersecurity red teaming gained significant momentum as organizations recognized the need for more comprehensive security testing. The Central Intelligence Agency created a new unit called “Red Cell”, and red teaming became increasingly common in various government agencies to model responses to asymmetric threats, including cyber attacks (Wikipedia, 2025). This period marked the transition from isolated pentesting to more holistic security assessments that incorporated physical security, social engineering, and other non-technical aspects.
@@
Modern cybersecurity red teaming encompasses several key methodologies working in concert: technical assessments that test digital defenses through vulnerability scanning, exploitation, and lateral movement (IBM, 2025); physical security testing that evaluates access controls for facilities (Sapphire, 2024); social engineering that targets the human element through phishing and impersonation (Tomkiel, 2024); and extended red team operations designed to achieve specific objectives while testing detection and response capabilities (Redscan, 2024). These approaches are codified in frameworks such as NIST Special Publication 800-53, which includes specific controls for red team exercises designed to “simulate attempts by adversaries to compromise organizational information systems” and “provide comprehensive assessments that reflect real-world conditions” (CSF Tools, 2021).
@@
The field continues to evolve with several advanced approaches. Continuous Automated Red Teaming (CART) uses automation to assess security posture in real-time rather than through periodic manual assessments (IBM, 2025). Adversary Emulation models tactics after specific threat actors that might target the organization, guided by frameworks like MITRE ATT&CK (CISA, 2024). Purple Teaming fosters collaboration between red and blue teams to identify vulnerabilities and improve response strategies (DePalma, 2023). Integrated IT-OT Assessments expand scope to include industrial control systems and critical infrastructure (Frumento, 2021). AI-Enhanced Red Teaming that incorporates AI to improve the effectiveness of assessments (Kirvan, 2024). And finally, specialized services from organizations like CISA provide comprehensive evaluations for critical infrastructure sectors and government agencies (CISA, 2024).
@@ 2.3 Red Teaming AI
A working definition of the very new concept of AI red teaming is that it is structured testing to identify flaws and vulnerabilities in AI systems, typically conducted in controlled environments with developer collaboration. For Large Language Models (LLMs) specifically, red teaming is defined as “a process where participants interact with the LLM under test to help uncover incorrect or harmful behaviors” (Romero-Arjona et al., 2025). LLM development companies have implemented various approaches to red teaming, ranging from comprehensive security assessments to narrower evaluations focused on specific genAI features (Bullwinkel et al., 2025). In the last two years, such approaches have become recognized as essential for assessing the risks of AI models and systems (Ahmad et al., 2025).
Despite its growing popularity, researchers have identified significant challenges with current AI red teaming practices. The field lacks consensus on the scope, structure, and assessment criteria for AI red teaming (Feffer et al., 2024), raising concerns that red teaming may sometimes function more as performative activity than substantive risk mitigation. Many current approaches focus too narrowly on the models themselves, neglecting how vulnerabilities might manifest in production systems where AI models are a part of broader systems (Bullwinkel et al., 2025). Additionally, most AI experts fail to consider insider risks (Martin and Mercer, 2025), and most testing processes remain limited to English-language evaluations (Romero-Arjona et al., 2025).
The process of red teaming itself presents challenges for participants. They can experience negative psychological impacts from interaction with harmful content and adversarial thinking, which can lead to decreased productivity or psychological harm (Ahmad et al., 2025). Many red teamers also lack training in crucial disciplines outside of their technical expertise, with most employee red teamers having limited proficiency in linguistic, sociocultural, historical, legal, or ethical domains (Gillespie et al., 2025).
As the field matures, researchers increasingly recognize that red teaming alone cannot solve all the challenges in AI risk assessment (Ahmad et al., 2025). More effective approaches such as violet teaming are emerging that integrate complementary methods, combining red teaming’s risk identification capabilities with the solution-focused responses of blue teaming (Titus and Russell, 2023). Future success will likely depend on embracing greater diversity and focusing more on practical red teaming efforts that address real-world attacks, which are often less sophisticated than those presented in academic papers (Rawat et al., 2024).
@@ 2.4 What’s Missing?
As such, the state-of-the-art of AI red teaming over-indexes on model-specific behaviors rather than how these models work within AI applications, interact with broader social contexts, and affect the external world through their outputs. Current AI red teaming efforts tend to focus narrowly on technical vulnerabilities while overlooking broader sociotechnical considerations. The public interest dimension of red teaming remains underdeveloped as well (Singh et al., 2025).
To mitigate these flaws, AI red teams need to embrace cybersecurity’s decades of experience with security testing and reporting. A large proportion of the testing done in the name of red teaming AI systems can be structured and automated through frameworks like MITRE ATT&CK for understanding adversarial behaviors, prioritizing vulnerabilities and coordinating defensive responses (MITRE, 2023), and established practices for continuous monitoring, automated testing, and incident response (NIST, 2020). In comparison, effective red teaming benefits from “an alchemist mindset” that extends beyond purely technical approaches (Inie et al., 2025). Successful cyber red team engagements typically involve creating diverse and highly realistic scenarios that produce actionable insights to aid proactive mitigation (Yulianto et al., 2025)—principles that are often missing from AI red teaming.
@@ 3.1 Macro-level (System) Red Teaming
Just like technical debt in ML systems can arise from system components other than code (Sculley et al., 2015), ML (and AI) failures can stem from decisions made long before the first line of code is written (Figure 1).
Figure 1: Technical and nontechnical components of a ML system, with components having potential trust considerations marked by blue circles. Reproduced with permission from Pruksachatkun et al. (2023).
@@ 3.1.1 Inception
The inception stage is critical, given this is where stakeholders first envision an AI solution to address a particular challenge. During this stage, the job of macro-level red teaming is to challenge fundamental assumptions about problem framing and solution appropriateness. Red teams at this stage should ask probing questions: Is AI actually necessary to solve this problem, or are we succumbing to “AI solutionism”? What are the stakeholder motivations driving this project, and how might different constituencies be affected? What adjacent problems might we inadvertently create? How might adversaries exploit or manipulate this system once deployed? (TRADCOC G2, 2015)
@@
Critically, they should help everyone be on the same page to answer the question “What does good look like?” To this end, stakeholders must articulate not just what they want the system to do, but what they absolutely cannot allow it to do. This negative space definition often reveals assumptions and constraints that were not initially apparent (Singh et al., 2025).